### v1

@@ Appendix B Deep imaging data reduction
Each raw frame was first corrected for a master bias (median‐stack of 21 bias frames), saturated pixels were masked, and the flat-field correction was performed using twilight flats. A comprehensive sky mask using aggresive Gnuastro’s NoiseChisel (Akhlaghi & Ichikawa, 2015) segmentation was applied, and the sky level was measured via a sigma‐clipped median of the unmasked pixels and subtracted, ensuring removal of background without oversubtracting the faint coma.
An initial WCS solution was obtained using Astrometry.net (Lang et al., 2010) against Gaia DR3, refined through three SCAMP (Bertin, 2006) iterations. Each image was then cropped to a 3’3’ region centered on the comet, using coordinates provided by JPL Horizons. To avoid potential astrometric uncertainties, the center was re-calculated from the object’s brightness centroid. Cutout images were reastrometrized and resampled onto a common grid with SWarp (Lanczos3 method, Bertin et al., 2002).
After calibration and sky subtraction, we examined all 326 frames for residual contamination (e.g., faint field stars or tracking artifacts) and excluded 99 images with any potential contamination to the comet signal. The remaining 227 frames, totaling 3.15 hours of integration, were combined using an inverse variance–weighted mean, where each image’s weight was set by the reciprocal of its background variance. In this combination, Noisechisel was used for masking every source but the comet and a 3 clipping was applied to reject outliers.
@@ Appendix C Gas production rate computation
To study the presence of typical cometary emission lines/bands of CN, C2, C3 or NH2 usually detected in the visible range in active comets (see, e.g., Cochran et al. 2012), we first obtained the continuum-subtracted spectrum of 3I/ATLAS (see Fig.5) using a spectrum of the Sun downloaded from the CALSPEC compilation (Bohlin et al., 2014) as we did in de León et al. (2020). We first noticed that there is no evidence of any gas emission in our spectrum. As the CN (0–0) emission band at 3880 Å use to be the most conspicuous one in comets, we computed the upper limit of the detectable CN flux within the SNR of our spectrum in the 3830–3905 Å range (see Fig.5 inset plot) and derived the upper limit of the CN production rate (Q(CN)lim) as follows: i) we first computed the of the flux in the continuum-subtracted spectrum in the 3830–3905 Å spectral range and obtained a value of ; ii) assuming a detection limit of 3- as ; iii) we finally converted into column density using the factor from Schleicher (2010), scaled to both the heliocentric distance and the velocity of the comet. To compute the gas production rate, we assumed the Haser modeling with the outflow velocity scaled with (= km s-1), customary values for the daughter velocity =1 km s-1 , and scale lengths given in A’Hearn et al. (1995). For the corresponding set of parameters in the Haser modeling, we produced theoretical column density profiles for CN varying the production rate until the best match between observations and theoretical predictions was achieved. We obtained mol s-1 .
@@ Appendix D Period-finding method
All raw TTT frames were first corrected for bias and twilight flat‐field variations using standard procedures. Point‐source photometry was then performed with a GPU‐accelerated custom pipeline that employs PSF‐matched convolutions for source detection and adaptive, S/N‐driven aperture photometry with PSF‐based aperture corrections. For each exposure, we matched field stars to the Pan-STARRS PS1-DR1 catalog (Chambers et al., 2016) to derive a photometric zero-point, which was used to calibrate the comet and control-star magnitudes in the resulting time series. This pipeline has been validated in previous works (Popescu et al., 2025) and will be released as open-source software (Alarcon et al., in prep.).
To determine the rotation period of 3I/ATLAS from the TTT time‐series photometry, we began by computing a generalized Lomb–Scargle periodogram (Press & Rybicki, 1989). This method is well suited for unevenly sampled data and quantifies the significance of sinusoidal signals by fitting 2-order sine curves at trial frequencies while properly normalizing for observational errors. Applying it to the combined ‐band light curve revealed a peak at h. To corroborate this result without assuming purely sinusoidal variability, we employed the Phase Dispersion Minimization (PDM) technique, which searches for the period that minimizes the variance of data points within phase bins. Our PDM statistic exhibited a global minimum at a similar period identified by the Lomb–Scargle analysis. We then fitted a Gaussian function to the PDM curve in the vicinity of this minimum (see Fig. 6), yielding a refined rotation period of h. The quoted uncertainty corresponds to the standard deviation of the Gaussian fit and encompasses both measurement noise and sampling effects. We adopt this value as our best estimate in the main text.
@@
Appendix E Orbit data
The orbital solution in Table 2 is currently based on 484 observations with an observational time span of 53 days, it is referred to epoch JD 2460860.5 TDB, which is the origin of time in the calculations. It was retrieved from Jet Propulsion Laboratory’s (JPL) Small-Body Database (SBDB)222https://ssd.jpl.nasa.gov/tools/sbdb_lookup.html#/ provided by the Solar System Dynamics Group (SSDG, Giorgini 2011, 2015).333https://ssd.jpl.nasa.gov/ Its current path is hyperbolic with a statistical significance (barycentric) above 645. No known processes, close encounters with known planets included, can explain such a large eccentricity if we were to assume that 3I/ATLAS has an origin in the solar system; an extrasolar origin emerges as the single other alternative.
@@ Appendix E Orbit data
| Perihelion distance, (au) | = | 1.35800.0011 | 1.3633 |
| Eccentricity, | = | 6.1510.008 | 6.163 |
| Inclination, (°) | = | 175.11370.0006 | 175.1276 |
| Longitude of the ascending node, (°) | = | 322.1700.011 | 322.259 |
| Argument of perihelion, (°) | = | 127.9950.011 | 128.073 |
| Mean anomaly, (°) | = | -855.30.9 | -854.7 |

### v2

@@ Appendix B Deep imaging data reduction
Each raw frame was first corrected for a master bias (median‐stack of 21 bias frames), saturated pixels were masked, and the flat-field correction was performed using twilight flats. A comprehensive sky mask using aggresive Gnuastro’s NoiseChisel (Akhlaghi & Ichikawa 2015) segmentation was applied, and the sky level was measured via a sigma‐clipped median of the unmasked pixels and subtracted, ensuring removal of background without oversubtracting the faint coma.
An initial WCS solution was obtained using Astrometry.net (Lang et al. 2010) against Gaia DR3, refined through three SCAMP (Bertin 2006) iterations. Each image was then cropped to a 3’3’ region centered on the comet, using coordinates provided by JPL Horizons. To avoid potential astrometric uncertainties, the center was re-calculated from the object’s brightness centroid. Cutout images were reastrometrized and resampled onto a common grid with SWarp (Lanczos3 method, Bertin et al. 2002). Photometric calibration was performed with reference to the Pan-STARRS DR1 survey (Chambers et al. 2016). We cross-matched the TTT field with Gaia DR3 (Gaia Collaboration et al. 2023) to isolate bona fide stars, selected point sources by their FWHM, and performed aperture photometry with radii that enclose the PSF wings on both image sets, only retaining unsaturated stars with high signal-to-noise. Fluxes from both data sets were referenced to Gaia BP/RP spectra to remove the small zero‑point offset between surveys, and a colour correction based on the index and the respective filter–transmission curves accounted for filter‑response mismatches. The final scale factor places the TTT images on the AB system with a global zero‑point of 22.5 mag. this factor was applied to every sky‑subtracted frame before stacking, thereby converting pixel values from ADU to physical flux units.
After calibration, we examined all 326 frames for residual contamination (e.g., faint field stars or tracking artifacts) and excluded 99 images with any potential contamination to the comet signal. The remaining 227 frames, totaling 3.15 hours of integration, were combined using an inverse variance–weighted mean, where each image’s weight was set by the reciprocal of its background variance. In this combination, Noisechisel was used for masking every source but the comet and a 3 clipping was applied to reject outliers.
@@ Appendix C Gas production rate computation
To study the presence of typical cometary emission lines/bands of CN, C2, C3 or NH2 usually detected in the visible range in active comets (see, e.g., Cochran et al. 2012), we first obtained the continuum-subtracted spectrum of 3I/ATLAS (see Fig.4) using a spectrum of the Sun downloaded from the CALSPEC compilation (Bohlin et al. 2014) as we did in de León et al. (2020). Notice that the noisier part of the spectrum falls below 3800 Å and thus it does not affect our computations. We first noticed that there is no evidence of any gas emission in our spectrum. As the CN (0–0) emission band at 3880 Å use to be the most conspicuous one in comets, we computed the upper limit of the detectable CN flux within the SNR of our spectrum in the 3830–3905 Å range (see Fig.4 inset plot) and derived the upper limit of the CN production rate (Q(CN)lim) as follows: i) we first computed the of the flux in the continuum-subtracted spectrum in the 3830–3905 Å spectral range and obtained a value of ; ii) assuming a detection limit of 3- as ; iii) we finally converted into column density using the factor from Schleicher (2010), scaled to both the heliocentric distance and the velocity of the comet. To compute the gas production rate, we assumed the Haser modeling with the outflow velocity scaled with (= km s-1), customary values for the daughter velocity =1 km s-1 , and scale lengths given in A’Hearn et al. (1995). We obtained mol s-1 .
@@ Appendix D Period-finding method
All raw TTT frames were first corrected for bias and twilight flat‐field variations using standard procedures. Point‐source photometry was then performed with a GPU‐accelerated custom pipeline that employs PSF‐matched convolutions for source detection and adaptive, S/N‐driven aperture photometry with PSF‐based aperture corrections. For each exposure, we matched field stars to the Pan-STARRS DR1 catalog (Chambers et al. 2016) to derive a photometric zero-point, which was used to calibrate the comet and control-star magnitudes in the resulting time series. This pipeline has been validated in previous works (Popescu et al. 2025) and will be released as open-source software (Alarcon et al., in prep.).
To determine the rotation period of 3I/ATLAS from the TTT time‐series photometry, we began by computing a generalized Lomb–Scargle periodogram (Press & Rybicki 1989). This method is well suited for unevenly sampled data and quantifies the significance of sinusoidal signals by fitting 2-order sine curves at trial frequencies while properly normalizing for observational errors. Applying it to the combined ‐band light curve revealed a peak at h. To corroborate this result without assuming purely sinusoidal variability, we employed the Phase Dispersion Minimization (PDM) technique, which searches for the period that minimizes the variance of data points within phase bins. Our PDM statistic exhibited a global minimum at a similar period identified by the Lomb–Scargle analysis. We then fitted a Gaussian function to the PDM curve in the vicinity of this minimum (see Fig. 5), yielding a refined rotation period of h. The quoted uncertainty corresponds to the standard deviation of the Gaussian fit and encompasses both measurement noise and sampling effects. We adopt this value as our best estimate in the main text.
@@
Appendix E Orbit data and post-perihelion evolution
The orbital solution in Table 2 is currently based on 762 observations with an observational time span of 60 days, it is referred to epoch JD 2460868.5 TDB, which is the origin of time in the calculations. It was retrieved from Jet Propulsion Laboratory’s (JPL) Small-Body Database (SBDB)333https://ssd.jpl.nasa.gov/tools/sbdb_lookup.html#/ provided by the Solar System Dynamics Group (SSDG, Giorgini 2011, 2015).444https://ssd.jpl.nasa.gov/ Its current path is hyperbolic with a statistical significance (barycentric) above 1718. No known processes, close encounters with known planets included, can explain such a large eccentricity if we were to assume that 3I/ATLAS has an origin in the solar system; an extrasolar origin emerges as the single other alternative.
@@ Appendix E Orbit data and post-perihelion evolution
Perihelion distance, (au) | = | 1.35700.0005 | 1.3623 |
Eccentricity, | = | 6.1440.003 | 6.156 |
Inclination, (°) | = | 175.11350.0002 | 175.1275 |
Longitude of the ascending node, (°) | = | 322.1640.005 | 322.254 |
Argument of perihelion, (°) | = | 128.0070.004 | 128.086 |
Mean anomaly, (°) | = | -796.30.3 | -795.8 |
Answering the question of where 3I/ATLAS is headed next after its fly-by with the Sun requires the analysis of a similar set of direct -body simulations (also with MCCM), but forward in time. At 1.78230.0003 pc from the Sun and 3104 yr into the future, this object will be receding from us at km s-1 towards (apex) , (9523006, 1980390.0005) in the constellation of Gemini with Galactic coordinates , , and ecliptic coordinates , . The heliocentric components of its Galactic velocity will be km s-1.